OpenStudy (itseasy2346):

how do you solve (6+4)^2+(9/3-1)?

OpenStudy (anonymous):

it equals 102

OpenStudy (kendricklamar2014):

@willowwest17 It clearly says solve not give a direct answer

OpenStudy (kendricklamar2014):

Here is how you solve this problem: \[(6+4)^2+(\frac{ 9 }{ 3 }-1)\] \[=10^2+ \frac{ 9 }{ 3 } −1\] \[=100+ \frac{ 9 }{ 3 } −1\] \[=100+3−1\] \[= 100 + 2\] \[= 102\]

OpenStudy (itseasy2346):

Thank you so much. :)

OpenStudy (kendricklamar2014):

Your welcome :)

OpenStudy (itseasy2346):

but how did you get 100+ 3-1

OpenStudy (kendricklamar2014):

\[\frac{ 9 }{ 3 } = 3\]

OpenStudy (itseasy2346):

oh ya right thanks.

OpenStudy (kendricklamar2014):

No problem

OpenStudy (anonymous):

lol wow. listen I plugged it in omy graphing calc. sorry I didn't do all the work for you. but I would realy appreciate if you didn't disrespect me. kaythanks!(:

OpenStudy (itseasy2346):

lol no its fine

OpenStudy (itseasy2346):

I'd never do that

OpenStudy (kendricklamar2014):

Well you dont need a calculator to solve this problem. All you need is mental math, or a pencil and paper. Im not disrespecting you, im just telling you to show steps so @itseasy2346 can learn how to solve this problem.

OpenStudy (itseasy2346):

I'm not so mean.

OpenStudy (itseasy2346):

and ya @KendrickLamar2014 is right.

OpenStudy (anonymous):

I understand that, but if you show steps this person will never know. I give the answer then they figure it out. I am taking dual credit classes at the University of Oregon, this is how a math teacher does it. its called self-sufficient work. it will help. u don't just give ur work. u give a lead then THEY show their work. anyways, im done. bye

OpenStudy (itseasy2346):

lol I'm in elementary in 8th grade and the teachers are really strict.

OpenStudy (itseasy2346):

so i just needed the work

OpenStudy (kendricklamar2014):

Well this isn't the University of Oregon, this is OpenStudy. Here you shows steps so the person understand. If your teacher does it another way, I understand. But here its different.

OpenStudy (itseasy2346):

Thanks you guys anyway.

OpenStudy (anonymous):

.... 8th grade? nvm.